OpenStudy (anonymous):

soup can r=6.6cm h=10cm

OpenStudy (anonymous):

Do you want surface area, volume, or what?

OpenStudy (anonymous):

Well, visualizing the surface area, it's 2 circles (the top and bottom) plus the side, which is essentially a rectangle with a length equal to the circumference of the top and bottom. Circle area = pi*r^2 Circumference = 2*pi*r S = 2(circle area) + h(circumference) S = 2(pi*6.6^2) + 10(2*pi*6.6) You can simplify that. For volume, it's (arguably) easier. It's one circle (the top/bottom), and you drag it out to the height of the whole thing. So it's the circle area times the height. V = h(circle area) V = 10(pi*6.6^2) You can simplify that.